In a basketball tournament each game last1 1/4 hours. In a 24 hour period what is the maximum number of games that can be played
Answer by stanbon(75887)

In a basketball tournament each game last1 1/4 hours. In a 24 hour period what is the maximum number of games that can be played
--------------
Ans: 24/(1 1/4) = 24/(5/4) = 24(4/5) = 19.2
Round down to 19 games
----
cheers,
Stan H.